OpenStudy (anonymous):

What is the inverse of the equation y = 5x - 1?

OpenStudy (linn99123):

y = 5x - 1 x = 5y - 1 f^-1 = 5y - 1 i think?

OpenStudy (anonymous):

hmm no

OpenStudy (anonymous):

\[y=5x-1\] means multiply by 5, subtract 1 do the inverse operations in the opposite order like if you were going to solve \[5x-1=12\] what would you do?

OpenStudy (anonymous):

divide both sides by 5x

OpenStudy (anonymous):

no, not if you want to solve \[5x-1=12\] the first step would not be to divide

OpenStudy (anonymous):

5x+11?

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

suppose you had an elementary algebra problem that said "solve for \(x\) if \(5x-1=12\) it would take two steps to solve this problem for \(x\) any idea how you would do this?